|a|=(2*n+1);
aba^(-1)=b^(-1);
b^2=?
1
Expert's answer
2020-11-02T19:07:15-0500

Given that,

"|a|=2^{n+1} \ and \ aba^{-1}=b^{-1}"

we have to find "b^2=?"

since "aba^{-1}=b^{-1} \implies (aba^{-1})^{-1}=(b^{-1})^{-1}"

"\implies ab^{-1}a^{-1}=b"

"\implies (ab^{-1}a^{-1})^2=b^2"

"\implies b^2=ab^{-1}a^{-1}ab^{-1}a^{-1}" "=ab^{-2}a^{-1}"